OpenStudy (anonymous):

A function has the form y(x,t)= e^(-1-(t+x)^2). a) Fixing the time at t=0, t=1, t=2, draw y as a function of x. b) Is this a wave? Determine the wave number k and circular frequency . c) What is the propagation velocity? Is it positive or negative? d) In what direction is it propagating (+x or –x)?

OpenStudy (anonymous):

@saifoo.khan @hartnn @dpaInc @JamesJ @cwrw238 @robtobey @radar

OpenStudy (anonymous):

thank you, i need help with b and c, I drew (a), and d I think the answer is that it is not moving

OpenStudy (anonymous):

You can say this is a wave because the function has the argument form \[x \pm v\] so x and t are combined in a linear fashion. You could also actually see if this is a solution of the partial differential equation of a wave but either argument is valid. so the argument of the function is called the phase which in this case is \[\phi =-1-(t+x)^2)\] so you are looking for the change of position with respect to time so you get \[0 =-(2tdt+(xdt+tdx)+2xdx)\] since the you are assuming the phase is constant or x and t are such that the phase is constant so get dt on one side and dx on the other \[2tdt+xdt = -tdx-2xdx\] factor out dx and dt \[dt(2t+x) = (-t-2x)dx\] finally isolate dx/dt \[dx/dt =-(2t+x)/(t+2x)\] would be a function for the velocity. So in this case it is moving in the negative x-direction as that whole function is negative.

OpenStudy (anonymous):

yes but still I do not get what would be the wave number k and w the angular frequency in this case? would they be as follow, k=1rad/m, w=1rad/s

OpenStudy (anonymous):

@saifoo.khan @hartnn @dpaInc @JamesJ @cwrw238 @robtobey @radar